# Multi-objective Aerial IRS-assisted ISAC Optimization via Generative AI-enhanced Deep Reinforcement Learning

### v1

@@ VI-C3 Impact of Maximum Power $P_{max}$
Notably, individually optimized objectives under communication-only and sensing-only methods outperform the corresponding objectives in the jointly optimized situation, which confirms that a significant trade-off relationship exists in the optimization objectives in the formulated CSEMOP. Moreover, in the communication-only method, the target echo rate approaches zero because the aerial ISR is deployed closer to the user cluster area and the passive beamforming of the IRS is fully aligned with the user, resulting in a significant attenuation of the sensing performance. In contrast, the user communication rate in the sensing-only method does not approach zero, likely due to the direct communication link between the BS and the users, which maintains a certain level of communication capability even under the sensing-priority condition. In addition, the random method performs poorly in all performance metrics due to the lack of systematic and theoretical basis for setting optimization variables by using a random strategy.
@@
VI-C4 UAV Trajectory Analysis
The schematic of the UAV trajectory under the spatial distribution of the BS, users, and target is shown in the Fig. 6. Initially, the ground location of aerial IRS is $(0\ \text{m},300\ \text{m})$. To efficiently perform the communication and sensing tasks, the aerial IRS flies toward the clustered region of the users, target and BS, and then hovers at its final location. Moreover, to validate the robustness of the proposed method, we conducted a repeated experiment with the initial location of the UAV set to $(300\ \text{m},0\ \text{m})$. The final hovering location is similar to that of the first simulation, further confirming the consistency and stability of the GDMDDPG method.
VII Conclusion
In this paper, we investigated an aerial IRS-assisted ISAC system, where a BS simultaneously provides communication and sensing services to multiple users and a target. Moreover, we formulated a CSEMOP to maximize the communication rate, echo rate while minimizing the UAV propulsion energy consumption by jointly optimizing BS beamforming matrix, IRS phase shifts, the flight velocity and angle of the UAV. Moreover, we proposed the GDMDDPG method, which improves the decision quality and learning efficiency by integrating the diffusion model, the noise perturbation mechanism, and the RPER sampling mechanism. Simulation results confirm that the proposed GDMDDPG method effectively accomplishes simultaneous optimization of communication, sensing, and energy efficiency performance, outperforming other comparison methods. In addition, GDMDDPG exhibits significant advantages in terms of robustness.

### v2

@@ 6.4.1 Impact of Maximum Transmit Power
Fig. 6 shows the trends of the three optimization objectives in CSEMOP with the variation of the BS maximum transmit power $P_{\rm max}$. The results demonstrate that as $P_{\rm max}$ increases, the communication rate and target sensing rate of DDPG, TD3, and the proposed GDMDDPG algorithms all show an upward trend. This is because the increasing $P_{\rm max}$ enhances signal strength, which improves both communication and sensing performance. Moreover, the GDMDDPG algorithm achieves the lowest UAV energy consumption, demonstrating its effectiveness in trajectory design. In addition, in the communication-only method, the target sensing rate approaches zero because the low-altitude ISR is deployed closer to the user cluster area and the passive beamforming of the IRS is fully aligned with the user, resulting in a significant attenuation of the sensing performance. In contrast, the user communication rate in the sensing-only method does not approach zero, likely due to the direct communication link between the BS and the users, which maintains a certain level of communication capability even under the sensing-priority condition. Furthermore, the random method performs poorly in all performance metrics due to the lack of systematic and theoretical basis for setting optimization variables by using a random strategy.
6.4.2 Impact of UAV Flight Height
Fig. 7 shows the average communication rate, sensing rate, and UAV energy consumption under different UAV flight heights. As can be seen, the average communication rate and sensing rate show a downtrend as the UAV flight height increases. This phenomenon occurs because the higher UAV flight height leads to higher path loss, which reduces channel gain and consequently impacts communication and sensing performance. Moreover, we find that the target sensing rate decreases more significantly than the user communication rate. This is primarily because direct communication links between the BS and users remain relatively unaffected by changes in UAV flight height. In addition, it can be observed that the proposed GDMDDPG algorithm outperforms the original DDPG algorithm, TD3 algorithm, and random strategy in terms of average communication rate, sensing rate, and UAV energy consumption. The result indicates that the proposed algorithm maintains effectiveness across different UAV flight heights, which is mainly due to the integrated diffusion model in enhancing agent comprehension of environmental states and decision quality.
Figure 8: The impact of antenna number on different optimization objectives. (a) Average communication rate. (b) Average sensing rate. (c) Average UAV energy consumption.
6.4.3 Impact of Antenna Number
Fig. 8 shows the optimization objective values under different numbers of antennas. As can be seen, the communication rate and sensing rate achieved by the GDMDDPG algorithm show an upward trend with increasing number of antennas. This is because more antennas enable for more accurate beamforming, which is beneficial in improving rate performance. Moreover, we can observe that when the number of antennas is 5, the communication rate and sensing rate obtained by the DDPG and TD3 algorithms decrease. This phenomenon occurs because the increasing number of antennas expands the action space dimension, making it difficult for DDPG and TD3 to find better solutions. Meanwhile, the above results indicate that the proposed GDMDDPG algorithm can still maintain outstanding exploration while the dimension of the action space increases.
@@ 6.4.3 Impact of Antenna Number
6.5 UAV Trajectory Analysis
The schematic of the UAV trajectory under the spatial distribution of the BS, users, and target is shown in Fig. 9. Initially, the location of UAV is $(0\ \text{m},300\ \text{m},40\ \text{m})$. To efficiently perform the communication and sensing tasks, the UAV flies toward the clustered region of the users, target and BS, and then hovers at its final location. Moreover, to validate the robustness of the proposed algorithm, we conduct a repeated simulation with the initial location of the UAV set to $(300\ \text{m},0\ \text{m},40\ \text{m})$. The final hovering location is similar to that of the first simulation, further confirming the consistency and stability of the GDMDDPG algorithm.
7 Conclusion
In this paper, we have investigated a low-altitude IRS-assisted ISAC system, where a BS simultaneously provides communication and sensing services to multiple users and a target. Moreover, we have formulated a CSEMOP to maximize the communication rate, sensing rate while minimizing the UAV propulsion energy consumption by jointly optimizing BS beamforming matrix, IRS phase shifts, the flight velocity and angle of the UAV. Moreover, we have proposed the GDMDDPG algorithm, which improves the decision quality and learning efficiency by integrating the diffusion model, the noise perturbation mechanism, and the RPER sampling mechanism. Simulation results have confirmed that the proposed GDMDDPG algorithm effectively accomplishes simultaneous optimization of communication, sensing, and energy efficiency performance, outperforming other comparison methods. In addition, GDMDDPG exhibits significant advantages in terms of robustness.